Solve by using the elimination method.
2r-9s=-10
9r+2s=40
What is the solution of the system ? Select the correct choice below and, if necessary, fill in the answer box to complete your choice.
A) the solution is_______(type an ordered pair. use integers or fractions for any numbers in the expression.)
B)There are infinitely number of solutions.
C) There is no solution.
Please help me with this mess!!!!!LOL

Answer by richwmiller(17219)

A) r=4 s=2
(4,2)
2r-9s=-10
9r+2s=40
Multiply equation #1 by 2
2*(2r-9s=-10)
Multiply equation #2 by 9
9*(9r+2s=40)
Add them
s is gone
Solve for r